# Integrating Large Language Models into a Tri-Modal Architecture for Automated Depression Classification

### v2

@@ Abstract
Major Depressive Disorder (MDD) is a pervasive mental health condition that affects 300 million people worldwide. This work presents a novel, BiLSTM-based tri-modal model-level fusion architecture for the binary classification of depression from clinical interview recordings. The proposed architecture incorporates Mel Frequency Cepstral Coefficients, Facial Action Units, and uses a two-shot learning based GPT-4 model to process text data. This is the first work to incorporate large language models into a multi-modal architecture for this task. It achieves impressive results on the DAIC-WOZ AVEC 2016 Challenge cross-validation split and Leave-One-Subject-Out cross-validation split, surpassing all baseline models and multiple state-of-the-art models. In Leave-One-Subject-Out testing, it achieves an accuracy of 91.01%, an F1-Score of 85.95%, a precision of 80%, and a recall of 92.86%.
@@ V Results
    The model was evaluated using the DAIC-WOZ AVEC 2016 Challenge cross-validation split to allow comparability with models from other studies.
@@
Figure 7: Confusion matrices of proposed architecture’s performance on AVEC 2016 cross-validation split (left) and LOSOCV (right).
TABLE VI: Comparison of proposed architecture with baseline binary depression classification models tested on the AVEC 2016 cross-validation split.
@@
TABLE VII: Comparison of proposed architecture with state-of-the-art binary depression classification models tested on the AVEC 2016 cross-validation split. A dash signifies that the metric was not reported in the paper. (*Evaluation done with LOSOCV)
@@ D ND D ND
The proposed model outperforms all baseline models and multiple state-of-the-art models which were evaluated using a regular cross-validation split as well as LOSOCV. It performs well on both the depressed and non-depressed classes, with accuracies of 92.9% and 90.2% on LOSOCV for the depressed and non-depressed classes, respectively. The large difference in the accuracy between the depressed and non-depressed classes (92.9% and 81.81%, respectively) on the AVEC 2016 cross-validation split can be attributed to the extremely small sample size of just 14 for the depressed class. We believe the accuracy of 92.9% for the depressed class on the AVEC 2016 Challenge cross-validation split is inflated as a result of this sample size.
@@ VII Conclusion
This paper presents a novel machine learning architecture for diagnosing MDD from clinical interview recordings. We explored the integration of large language models and audiovisual features into a tri-modal architecture for this task. The proposed architecture was optimized through automated hyper-parameter tuning techniques and evaluated using the AVEC 2016 Challenge cross-validation split and LOSOCV. We drew comparisons between the performance of the proposed architecture and previously developed models, finding that it outperforms all baseline models and multiple state-of-the-art models. The model was integrated into a locally hosted web application to emulate how it may be used in the real-world.

### v3

@@ Abstract
Major Depressive Disorder (MDD) is a pervasive mental health condition that affects 300 million people worldwide. This work presents a novel, BiLSTM-based tri-modal model-level fusion architecture for the binary classification of depression from clinical interview recordings. The proposed architecture incorporates Mel Frequency Cepstral Coefficients, Facial Action Units, and uses a two-shot learning based GPT-4 model to process text data. This is the first work to incorporate large language models into a multi-modal architecture for this task. It achieves impressive results on the DAIC-WOZ AVEC 2016 Challenge train/validation/test split and Leave-One-Subject-Out cross-validation split, surpassing all baseline models and multiple state-of-the-art models. In Leave-One-Subject-Out testing, it achieves an accuracy of 91.01%, an F1-Score of 85.95%, a precision of 80%, and a recall of 92.86%.
@@ V Results
    The model was evaluated using the DAIC-WOZ AVEC 2016 Challenge train/validation/test split to allow comparability with models from other studies.
@@
Figure 7: Confusion matrices of proposed architecture’s performance on AVEC 2016 train/validation/test split (left) and LOSOCV (right).
TABLE VI: Comparison of proposed architecture with baseline binary depression classification models tested on the AVEC 2016 train/validation/test split.
@@
TABLE VII: Comparison of proposed architecture with state-of-the-art binary depression classification models tested on the AVEC 2016 train/validation/test split. A dash signifies that the metric was not reported in the paper. (*Evaluation done with LOSOCV)
@@ D ND D ND
The proposed model outperforms all baseline models and multiple state-of-the-art models which were evaluated using no cross-validation split as well as LOSOCV. It performs well on both the depressed and non-depressed classes, with accuracies of 92.9% and 90.2% on LOSOCV for the depressed and non-depressed classes, respectively. The large difference in the accuracy between the depressed and non-depressed classes (92.9% and 81.81%, respectively) on the AVEC 2016 train/validation/test split can be attributed to the extremely small sample size of just 14 for the depressed class. We believe the accuracy of 92.9% for the depressed class on the AVEC 2016 Challenge cross-validation split is inflated as a result of this sample size.
@@ VII Conclusion
This paper presents a novel machine learning architecture for diagnosing MDD from clinical interview recordings. We explored the integration of large language models and audiovisual features into a tri-modal architecture for this task. The proposed architecture was optimized through automated hyper-parameter tuning techniques and evaluated using the AVEC 2016 Challenge train/validation/test split and LOSOCV. We drew comparisons between the performance of the proposed architecture and previously developed models, finding that it outperforms all baseline models and multiple state-of-the-art models. The model was integrated into a locally hosted web application to emulate how it may be used in the real-world.